# GVE-Leiden: Fast Leiden Algorithm for Community Detection in Shared Memory Setting

### v6

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents one of the most efficient parallel implementation of the Leiden algorithm, a high quality community detection method. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Leiden implementation, which we term as GVE-Leiden, outperforms the original Leiden, igraph Leiden, and NetworKit Leiden by 436×436$\times$436 ×, 104×104$\times$104 ×, and 8.2×8.2$\times$8.2 × respectively - achieving a processing rate of $403M$ edges/s on a $3.8B$ edge graph. Compared to GVE-Louvain, our parallel Louvain implementation, GVE-Leiden achieves a total elimination of disconnected communities, with only a $13\%$ increase in runtime. In addition, GVE-Leiden improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection is the problem of identifying subsets of vertices that exhibit higher connectivity among themselves than with the rest of the network. The identified communities are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community. These communities, also known as clusters, shed light on the organization and functionality of the network. It is an NP-hard problem with applications in topic discovery, protein annotation, recommendation systems, and targeted advertising (Gregory, 2010). The Louvain method (Blondel et al., 2008) is a popular heuristic-based approach for community detection. It employs a two-phase approach, comprising a local-moving phase and an aggregation phase, to iteratively optimize the modularity metric — a measure of community quality (Newman, 2006).
@@
In this paper, we present our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/leiden-communities-openmp. It incorporates several optimizations, including parallel prefix sums, preallocated Compressed Sparse Row (CSR) data structures for community vertex identification and super-vertex graph storage during aggregation, fast collision-free per-thread hash tables for the local-moving and aggregation phases, and prevention of unnecessary aggregations. Additionally, we employ a greedy refinement phase where vertices optimize for delta-modularity within their community bounds, yielding improved performance and quality compared to a randomized approach. Furthermore, we utilize established techniques such as asynchronous computation, OpenMP’s dynamic loop schedule, threshold-scaling optimization, and vertex pruning. To the best of our knowledge, our implementation is the most efficient implementation of Leiden algorithm on multicore CPUs to date. We conduct comprehensive comparisons with other state-of-the-art implementations, including multi-core, multi-node implementations, detailed in Table 1. Both direct and indirect comparisons are provided, with further details outlined in Sections 5.2 and A.3, respectively.
Table 1. Speedup of our multicore implementation of Leiden algorithm compared to other state-of-the-art implementations. Direct comparisons entail running the given implementation on our server, while indirect comparisons (marked with a *, explained in Section A.3) involve comparing results relative to a common reference.
Leiden implementation Published Our Speedup
Original Leiden (Traag et al., 2019) 2019 $\times2$
igraph Leiden (Csardi et al., 2006) 2006 50×50$\times$50 ×
NetworKit Leiden (Staudt et al., 2016) 2016 20×20$\times$20 ×
KatanaGraph Leiden (kat, 2023) 2023 $166\times^{*}$
ParLeiden-S (Hu et al., [n. d.]) 2023 $22\times^{*}$
ParLeiden-D (8 nodes) (Hu et al., [n. d.]) 2023 $18\times^{*}$
@@ 2. Related work
We now discuss a number of algorithmic improvements proposed for the Louvain method, that also apply to the Leiden algorithm. These include ordering of vertices based on importance (Aldabobi et al., 2022), attempting local move only on likely vertices (Ryu and Kim, 2016; Ozaki et al., 2016; Zhang et al., 2021; Shi et al., 2021), early pruning of non-promising candidates (Ryu and Kim, 2016; Halappanavar et al., 2017; Zhang et al., 2021; You et al., 2022), moving vertices to a random neighbor community (Traag, 2015), subnetwork refinement (Waltman and Eck, 2013; Traag et al., 2019), multilevel refinement (Rotta and Noack, 2011; Gach and Hao, 2014; Shi et al., 2021), threshold cycling (Ghosh et al., 2018), threshold scaling (Lu et al., 2015; Naim et al., 2017; Halappanavar et al., 2017), and early termination (Ghosh et al., 2018). A number of parallelization techniques have been also attempted for the Louvain method, that may also be applied to the Leiden algorithm. These include using adaptive parallel thread assignment (Fazlali et al., 2017; Naim et al., 2017; Sattar and Arifuzzaman, 2019; Mohammadi et al., 2020), parallelizing the costly first iteration (Wickramaarachchi et al., 2014), ordering vertices via graph coloring (Halappanavar et al., 2017), performing iterations asynchronously (Que et al., 2015; Shi et al., 2021), and using sort-reduce instead of hashing (Cheong et al., 2013).
A few open source implementations and software packages have been developed for community detection using Leiden algorithm. The original implementation of the Leiden algorithm (Traag et al., 2019), called libleidenalg, is written in C++ and has a Python interface called leidenalg. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It utilizes a hybrid with C++ kernels and a Python frontend. The package features a parallel implementation of the Leiden algorithm by Nguyen (Nguyen, [n. d.]) which uses global queues for vertex pruning, and vertex and community locking for updating communities. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in academic research, and includes an implementation of the Leiden algorithm.
@@ 4.1. Optimizations for Leiden algorithm
We extend our optimization techniques, originally designed for the Louvain method (Sahu, 2023), to the Leiden algorithm. Specifically, we implement an asynchronous version of the Leiden algorithm, allowing threads to operate independently on distinct sections of the graph. While this approach promotes faster convergence, it also introduces variability into the final result (Shi et al., 2021). To ensure efficient computations, we allocate a dedicated hashtable per thread. These hashtables serve two main purposes: they keep track of the delta-modularity associated with moving to each community connected to a vertex during the local-moving/refinement phases, and they record the total edge weight between super-vertices in the aggregation phase of the algorithm (Sahu, 2023).
Our optimizations include utilizing OpenMP’s dynamic loop scheduling, capping the number of iterations per pass at $20$, employing a tolerance drop rate of $10$ (threshold scaling), initiating with a tolerance of $0.01$, using an aggregation tolerance of $0.8$ to avoid performing aggregations of minimal utility, implementing flag-based vertex pruning (instead of a queue-based one (Nguyen, [n. d.])), utilizing parallel prefix sum, and using preallocated CSRs for identifying community vertices and storing the super-vertex graph during aggregation. Additionally, we employ fast collision-free per-thread hashtables, well separated in their memory addresses (Sahu, 2023).
@@ 4.2. Our optimized Leiden implementation
We now explain the implementation of GVE-Leiden in Algorithms 1, 2, 3, and 4. A flow diagram illustrating the first pass of GVE-Leiden is shown in Figure 3.
@@ 4.2.1. Main step of GVE-Leiden
The main step of GVE-Leiden (leiden() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, the refinement phase, and the aggregation phase. Here, the leiden() function accepts the input graph $G$, and returns the community membership $C$ of each vertex. In line 15, we first initialize the community membership $C$ for each vertex in $G$, and perform passes of the Leiden algorithm, limited to $MAX\_PASSES$ (lines 16-28). During each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 17).
Subsequently, in line 18, we perform the local-moving phase by invoking leidenMove(), which optimizes community assignments. Following this, we set the community bound of each vertex (for the refinement phase) as the community membership of each vertex just obtained, and reset the membership of each vertex, and the total weight of each community as singleton vertices in line 19. In line 20, the refinement phase is carried out by invoking leidenRefine(), which optimizes the community assignment of each vertex within its community bound. If the local-moving phase converges in a single iteration, global convergence is implied and we terminate the passes (line 21). Further, if the drop in the number of communities $|\Gamma|$ is marginal, we halt the algorithm at the current pass (line 23).
If convergence has not been achieved, we proceed to renumber communities (line 24), update top-level community memberships $C$ with dendrogram lookup (line 25), perform the aggregation phase by calling leidenAggregate(), and adjust the convergence threshold for subsequent passes, i.e., perform threshold scaling (line 28). The next pass commences in line 16. At the end of all passes, we perform a final update of the top-level community memberships $C$ with dendrogram lookup (line 29), and return the top-level community membership $C$ of each vertex in $G$.
@@
7:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
8:$H_{t}$: Collision-free per-thread hashtable
9:$l_{i}$, $l_{j}$: Number of iterations performed (per pass)
10:$l_{p}$: Number of passes performed
11:$\tau$: Per iteration tolerance
12:$\tau_{agg}$: Aggregation tolerance
13:
14:function leiden($G$)
15:     Vertex membership: $C\leftarrow[0..|V|)$ ; $G^{\prime}\leftarrow G$
@@
17:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow vertexWeights(G^{\prime})$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$
18:         $l_{i}\leftarrow leidenMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},\tau)$triangleright Alg. 2
19:         $C^{\prime}_{B}\leftarrow C^{\prime}$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$ ; $\Sigma^{\prime}\leftarrow K^{\prime}$
20:         $l_{j}\leftarrow leidenRefine(G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime},\tau)$triangleright Alg. 3
21:         if $l_{i}+l_{j}\leq 1$ then break $\triangleright$ Globally converged?
22:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
23:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
24:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
25:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
26:         $G^{\prime}\leftarrow leidenAggregate(G^{\prime},C^{\prime})$triangleright Alg. 4
27:         $C^{\prime}\leftarrow$ Map $C^{\prime}$ to $C^{\prime}_{B}$triangleright Use move-based membership
@@
29:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
30:     return $C$
@@
5:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
6:$H_{t}$: Collision-free per-thread hashtable
7:$l_{i}$: Number of iterations performed
8:$\tau$: Per iteration tolerance
9:
10:function leidenMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},\tau$)
11:     Mark all vertices in $G^{\prime}$ as unprocessed
@@
13:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
14:         for all unprocessed $i\in V^{\prime}$ in parallel do
15:              Mark $i$ as processed (prune)
16:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
18:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
19:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
20:              if $c^{*}=C^{\prime}[i]$ then continue
21:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomic
22:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
23:              Mark neighbors of $i$ as unprocessed
24:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
25:     return $l_{i}$
26:
27:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
28:     for all $(j,w)\in G^{\prime}.edges(i)$ do
29:         if not $self$ and $i=j$ then continue
30:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
31:     return $H_{t}$
@@
5:$G^{\prime}_{C^{\prime}}$: Community vertices (CSR)
6:$H_{t}$: Collision-free per-thread hashtable
7:$\tau$: Per iteration tolerance
8:
9:function leidenRefine($G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime},\tau$)
10:     for all $i\in V^{\prime}$ in parallel do
11:         $c\leftarrow C^{\prime}[i]$
12:         if $\Sigma^{\prime}[c]\neq K^{\prime}[i]$ then continue
13:         $H_{t}\leftarrow scanBounded(\{\},G^{\prime},C^{\prime}_{B},C^{\prime},i,false)$
@@
15:         $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$ within $C^{\prime}_{B}$
16:         $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
17:         if $c^{*}=c$ then continue
18:         if $atomicCAS(\Sigma^{\prime}[c],K^{\prime}[i],0)=K^{\prime}[i]$ then
19:              $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomically
20:              $C^{\prime}[i]\leftarrow c^{*}$
21:
22:function scanBounded($H_{t},G^{\prime},C^{\prime}_{B},C^{\prime},i,self$)
23:     for all $(j,w)\in G^{\prime}.edges(i)$ do
24:         if not $self$ and $i=j$ then continue
25:         if $C^{\prime}_{B}[i]\neq C^{\prime}_{B}[j]$ then continue
26:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
27:     return $H_{t}$
28:
29:function atomicCAS($pointer,old,new$)
@@
31:     if $pointer=old$ then $pointer\leftarrow new$ ; return $old$
32:     else return $pointer$
@@
19:         if degree of $c$ in $G^{\prime}_{C^{\prime}}=0$ then continue
20:         $H_{t}\leftarrow\{\}$
21:         for all $i\in G^{\prime}_{C^{\prime}}.edges(c)$ do
22:              $H_{t}\leftarrow scanCommunities(H,G^{\prime},C^{\prime},i,true)$
23:         for all $(d,w)\in H_{t}$ do
24:              Add edge $(c,d,w)$ to CSR $G^{\prime\prime}$ atomically
25:     return $G^{\prime\prime}$

### v7

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents one of the most efficient implementations of the Leiden algorithm, a high quality community detection method. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Leiden implementation, which we term as GVE-Leiden, outperforms the original Leiden, igraph Leiden, NetworKit Leiden, and cuGraph Leiden (running on NVIDIA A100 GPU) by 436×436$\times$436 ×, 104×104$\times$104 ×, 8.2×8.2$\times$8.2 ×, and 3.0×3.0$\times$3.0 × respectively — achieving a processing rate of $403M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Leiden improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection is the problem of identifying subsets of vertices that exhibit higher connectivity among themselves than with the rest of the network. The identified communities are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community. These communities, also known as clusters, shed light on the organization and functionality of the network. This problem with applications in topic discovery, protein annotation, recommendation systems, and targeted advertising (Gregory, 2010). One of the difficulties in the community detection problem is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008). The Louvain method (Blondel et al., 2008) is a popular heuristic-based approach for community detection. It employs a two-phase approach, comprising a local-moving phase and an aggregation phase, to iteratively optimize the modularity metric — a measure of community quality (Newman, 2006).
@@
In this report, we present our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/leiden-communities-openmp. It incorporates several optimizations, including parallel prefix sums, preallocated Compressed Sparse Row (CSR) data structures for community vertex identification and super-vertex graph storage during aggregation, fast collision-free per-thread hash tables for the local-moving and aggregation phases, and prevention of unnecessary aggregations — enabling it to run significantly faster than existing implementations. Additionally, we employ a greedy refinement phase where vertices optimize for delta-modularity within their community bounds, yielding improved performance and quality compared to a randomized approach. Furthermore, we utilize established techniques such as asynchronous computation, OpenMP’s dynamic loop schedule, threshold-scaling optimization, and vertex pruning. To the best of our knowledge, our implementation is the most efficient implementation of Leiden algorithm on multicore CPUs to date. We conduct comprehensive comparisons with other state-of-the-art implementations, including multi-core, GPU-based implementations, multi-node implementations, detailed in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.2, respectively.
Table 1. Speedup of our multicore implementation of Leiden algorithm compared to other state-of-the-art implementations. Direct comparisons entail running the given implementation on our servers, while indirect comparisons (marked with a *, explained in Section A.2) involve comparing results relative to a common reference.
Leiden implementation Parallelism Our Speedup
Original Leiden (Traag et al., 2019) Sequential $\times2$
igraph Leiden (Csardi et al., 2006) Sequential 50×50$\times$50 ×
NetworKit Leiden (Staudt et al., 2016) Parallel 20×20$\times$20 ×
cuGraph Leiden (Kang et al., 2023) Parallel (GPU) 3.0×3.0$\times$3.0 ×
ParLeiden-S (Hu et al., [n. d.]) Parallel $22\times^{*}$
ParLeiden-D (8 nodes) (Hu et al., [n. d.]) Multi-node $18\times^{*}$
@@ 2. Related work
We now discuss a number of algorithmic improvements proposed for the Louvain method, that also apply to the Leiden algorithm. These include ordering of vertices based on importance (Aldabobi et al., 2022), attempting local move only on likely vertices (Ryu and Kim, 2016; Zhang et al., 2021; Shi et al., 2021), early pruning of non-promising candidates (Ryu and Kim, 2016; Halappanavar et al., 2017; Zhang et al., 2021; You et al., 2022), moving vertices to a random neighbor community (Traag, 2015), subnetwork refinement (Waltman and Eck, 2013; Traag et al., 2019), multilevel refinement (Rotta and Noack, 2011; Gach and Hao, 2014; Shi et al., 2021), threshold cycling (Ghosh et al., 2018), threshold scaling (Lu et al., 2015; Naim et al., 2017; Halappanavar et al., 2017), and early termination (Ghosh et al., 2018). A number of parallelization techniques have been also attempted for the Louvain method, that may also be applied to the Leiden algorithm. These include using adaptive parallel thread assignment (Fazlali et al., 2017; Naim et al., 2017; Sattar and Arifuzzaman, 2019; Mohammadi et al., 2020), parallelizing the costly first iteration (Wickramaarachchi et al., 2014), ordering vertices via graph coloring (Halappanavar et al., 2017), performing iterations asynchronously (Shi et al., 2021), and using sort-reduce instead of hashing (Cheong et al., 2013).
A few open source implementations and software packages have been developed for community detection using Leiden algorithm. The original implementation of the Leiden algorithm (Traag et al., 2019), called libleidenalg, is written in C++ and has a Python interface called leidenalg. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It utilizes a hybrid with C++ kernels and a Python frontend. The package features a parallel implementation of the Leiden algorithm by Nguyen (Nguyen, [n. d.]) which uses global queues for vertex pruning, and vertex and community locking for updating communities. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in academic research, and includes an implementation of the Leiden algorithm. cuGraph (Kang et al., 2023) is a GPU-accelerated graph analytics library, part of the RAPIDS suite of data science and machine learning libraries. It leverages the computational power of NVIDIA GPUs to perform graph analytics tasks much faster than traditional CPU-based approaches. cuGraph’s core is implemented in C++ with CUDA and is used primarily through its Python interface, making it convenient for data scientists and developers.
@@ 4.1. Optimizations for Leiden algorithm
We extend optimization techniques, originally designed for the Louvain method (Sahu, 2023b), to the Leiden algorithm. Specifically, we implement an asynchronous version of the Leiden algorithm, allowing threads to operate independently on distinct sections of the graph. While this approach promotes faster convergence, it also introduces variability into the final result (Shi et al., 2021). To ensure efficient computations, we allocate a dedicated hashtable per thread. These hashtables serve two main purposes: they keep track of the delta-modularity associated with moving to each community connected to a vertex during the local-moving/refinement phases, and they record the total edge weight between super-vertices in the aggregation phase of the algorithm (Sahu, 2023b).
Our optimizations include using preallocated Compressed Sparse Row (CSR) data structures for identifying community vertices ($G^{\prime}_{C^{\prime}}$ in Algorithm 4) and storing the super-vertex graph ($G^{\prime\prime}$ in Algorithm 4) during aggregation, utilizing parallel prefix sum (lines 10-11, 16 in Algorithm 4), employing fast collision-free per-thread hashtables that are well separated in their memory addresses ($H_{t}$ in Algorithms 2, 3, and 4), and using an aggregation tolerance of $0.8$ to avoid performing aggregations of minimal utility (line 21 in Algorithm 4). Additionally, we implement flag-based vertex pruning (lines 10, 14, 22 in Algorithm 2) — instead of a queue-based one (Nguyen, [n. d.]), utilize OpenMP’s dynamic loop scheduling, cap the number of iterations per pass at $20$ (line 11 in Algorithm 2), employ a tolerance drop rate of $10$ (line 26 in Algorithm 1) — threshold scaling optimization, and initiate with a tolerance of $0.01$ (Sahu, 2023b).
@@
We also attempt two different variations of Parallel Leiden algorithm, one where the community labels of super-vertices (upon aggregation) is based on the local-moving phase (move-based), and the other where the community labels of super-vertices is based on the refinement phase (refine-based). Our observations indicate that both approaches have roughly the same runtime and modularity on average, as indicated by Figures 3 and 4. Accordingly, we stick to the move-based approach, which is the one recommended by Traag et al. (Traag et al., 2019). However, refine-based approach may be more suitable for the design of dynamic Leiden algorithm (for dynamic graphs).
@@ 4.2. Our optimized Leiden implementation
We now explain the implementation of GVE-Leiden in Algorithms 1, 2, 3, and 4. Our intention is to integrate GVE-Leiden into a forthcoming command-line graph processing tool named ”GVE”, which simply stands for Graph(Vertices, Edges), hence the name. GVE-Leiden operates with a time complexity of $O(KM)$, where $K$ is the total number of iterations performed, and a space complexity of $O(TN+M)$, where $T$ represents the number of threads, and $TN$ accounts for the collision-free hash tables $H_{t}$ used per thread. Figure 5 illustrates the first pass of GVE-Leiden.
@@ 4.2.1. Main step of GVE-Leiden
The main step of GVE-Leiden (leiden() function) is outlined in Algorithm 1. It encompasses initialization, the local-moving phase, the refinement phase, and the aggregation phase. Here, the leiden() function accepts the input graph $G$, and returns the community membership $C$ of each vertex. In line 13, we first initialize the community membership $C$ for each vertex in $G$, and perform passes of the Leiden algorithm, limited to $MAX\_PASSES$ (lines 14-26). During each pass, we initialize the total edge weight of each vertex $K^{\prime}$, the total edge weight of each community $\Sigma^{\prime}$, and the community membership $C^{\prime}$ of each vertex in the current graph $G^{\prime}$ (line 15).
Subsequently, in line 16, we perform the local-moving phase by invoking leidenMove(), which optimizes community assignments. Following this, we set the community bound of each vertex (for the refinement phase) as the community membership of each vertex just obtained, and reset the membership of each vertex, and the total weight of each community as singleton vertices in line 17. In line 18, the refinement phase is carried out by invoking leidenRefine(), which optimizes the community assignment of each vertex within its community bound. If the local-moving phase converges in a single iteration, global convergence is implied and we terminate the passes (line 19). Further, if the drop in the number of communities $|\Gamma|$ is marginal, we halt the algorithm at the current pass (line 21).
If convergence has not been achieved, we proceed to renumber communities (line 22), update top-level community memberships $C$ with dendrogram lookup (line 23), perform the aggregation phase by calling leidenAggregate(), and adjust the convergence threshold for subsequent passes, i.e., perform threshold scaling (line 26). The next pass commences in line 14. At the end of all passes, we perform a final update of the top-level community memberships $C$ with dendrogram lookup (line 27), and return the top-level community membership $C$ of each vertex in $G$.
@@
7:$l_{i}$, $l_{j}$: Number of iterations performed (per pass)
8:$l_{p}$: Number of passes performed
9:$\tau$: Per iteration tolerance
10:$\tau_{agg}$: Aggregation tolerance
11:
12:function leiden($G$)
13:     Vertex membership: $C\leftarrow[0..|V|)$ ; $G^{\prime}\leftarrow G$
@@
15:         $\Sigma^{\prime}\leftarrow K^{\prime}\leftarrow vertexWeights(G^{\prime})$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$
16:         $l_{i}\leftarrow leidenMove(G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},\tau)$triangleright Alg. 2
17:         $C^{\prime}_{B}\leftarrow C^{\prime}$ ; $C^{\prime}\leftarrow[0..|V^{\prime}|)$ ; $\Sigma^{\prime}\leftarrow K^{\prime}$
18:         $l_{j}\leftarrow leidenRefine(G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime})$triangleright Alg. 3
19:         if $l_{i}+l_{j}\leq 1$ then break $\triangleright$ Globally converged?
20:         $|\Gamma|,|\Gamma_{old}|\leftarrow$ Number of communities in $C$, $C^{\prime}$
21:         if $|\Gamma|/|\Gamma_{old}|>\tau_{agg}$ then break $\triangleright$ Low shrink?
22:         $C^{\prime}\leftarrow$ Renumber communities in $C^{\prime}$
23:         $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
24:         $G^{\prime}\leftarrow leidenAggregate(G^{\prime},C^{\prime})$triangleright Alg. 4
25:         $C^{\prime}\leftarrow$ Map $C^{\prime}$ to $C^{\prime}_{B}$triangleright Use move-based membership
@@
27:     $C\leftarrow$ Lookup dendrogram using $C$ to $C^{\prime}$
28:     return $C$
@@
5:$H_{t}$: Collision-free per-thread hashtable
6:$l_{i}$: Number of iterations performed
7:$\tau$: Per iteration tolerance
8:
9:function leidenMove($G^{\prime},C^{\prime},K^{\prime},\Sigma^{\prime},\tau$)
10:     Mark all vertices in $G^{\prime}$ as unprocessed
@@
12:         Total delta-modularity per iteration: $\Delta Q\leftarrow 0$
13:         for all unprocessed $i\in V^{\prime}$ in parallel do
14:              Mark $i$ as processed (prune)
15:              $H_{t}\leftarrow scanCommunities(\{\},G^{\prime},C^{\prime},i,false)$
@@
17:              $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$
18:              $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
19:              if $c^{*}=C^{\prime}[i]$ then continue
20:              $\Sigma^{\prime}[C^{\prime}[i]]-=K^{\prime}[i]$ ; $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomically
21:              $C^{\prime}[i]\leftarrow c^{*}$ ; $\Delta Q\leftarrow\Delta Q+\delta Q^{*}$
22:              Mark neighbors of $i$ as unprocessed
23:         if $\Delta Q\leq\tau$ then break $\triangleright$ Locally converged?
24:     return $l_{i}$
25:
26:function scanCommunities($H_{t},G^{\prime},C^{\prime},i,self$)
27:     for all $(j,w)\in G^{\prime}.edges(i)$ do
28:         if not $self$ and $i=j$ then continue
29:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
30:     return $H_{t}$
@@
5:$H_{t}$: Collision-free per-thread hashtable
6:
7:function leidenRefine($G^{\prime},C^{\prime}_{B},C^{\prime},K^{\prime},\Sigma^{\prime}$)
8:     for all $i\in V^{\prime}$ in parallel do
9:         $c\leftarrow C^{\prime}[i]$
10:         if $\Sigma^{\prime}[c]\neq K^{\prime}[i]$ then continue
11:         $H_{t}\leftarrow scanBounded(\{\},G^{\prime},C^{\prime}_{B},C^{\prime},i,false)$
@@
13:         $c^{*}\leftarrow$ Best community linked to $i$ in $G^{\prime}$ within $C^{\prime}_{B}$
14:         $\delta Q^{*}\leftarrow$ Delta-modularity of moving $i$ to $c^{*}$
15:         if $c^{*}=c$ then continue
16:         if $atomicCAS(\Sigma^{\prime}[c],K^{\prime}[i],0)=K^{\prime}[i]$ then
17:              $\Sigma^{\prime}[c^{*}]+=K^{\prime}[i]$ atomically ; $C^{\prime}[i]\leftarrow c^{*}$
18:
19:function scanBounded($H_{t},G^{\prime},C^{\prime}_{B},C^{\prime},i,self$)
20:     for all $(j,w)\in G^{\prime}.edges(i)$ do
21:         if not $self$ and $i=j$ then continue
22:         if $C^{\prime}_{B}[i]\neq C^{\prime}_{B}[j]$ then continue
23:         $H_{t}[C^{\prime}[j]]\leftarrow H_{t}[C^{\prime}[j]]+w$
24:     return $H_{t}$
25:
26:function atomicCAS($pointer,old,new$)
@@
28:     if $pointer=old$ then $pointer\leftarrow new$ ; return $old$
29:     else return $pointer$
@@
19:         $H_{t}\leftarrow\{\}$
20:         for all $i\in G^{\prime}_{C^{\prime}}.edges(c)$ do
21:              $H_{t}\leftarrow scanCommunities(H_{t},G^{\prime},C^{\prime},i,true)$
22:         for all $(d,w)\in H_{t}$ do
23:              Add edge $(c,d,w)$ to CSR $G^{\prime\prime}$ atomically
24:     return $G^{\prime\prime}$